# Uniqueness of free 2-periodicities of links

### v1

@@ 1. Introduction
Links in the real projective 3-space $\mathbb{R}\mathbb{P}^{3}$ (also known as projective links) have been investigated in various ways, often using diagrams, as in [16, 23, 24]. More generally, links in lens spaces have been considered [1, 11]. For a link $L$ in a lens space, the preimage (also called the lift) of $L$ by the universal covering map of the lens space is a freely periodic link in the 3-sphere $S^{3}$ according to Hartley [18]. Free periodicities of links in $S^{3}$ are quite restricted, especially for knots. For instance, Sakuma [26] and Boileau and Flapan [2] independently proved that a free periodicity of a prime knot in $S^{3}$ is unique. Several constraints for freely periodic knots in $S^{3}$ are known [5, 6, 12, 13, 18, 25]. Compatibility of free periodicities and alternating diagrams has also been considered [4, 14]. Furthermore, the finite group actions on knots in $S^{3}$ were classified in [7].
In this paper, we consider links in $\mathbb{R}\mathbb{P}^{3}$ and their preimages in $S^{3}$. Every link is assumed to be tame. Let $\pi\colon S^{3}\to\mathbb{R}\mathbb{P}^{3}$ denote the double covering map. Two links $L_{0}$ and $L_{1}$ in $\mathbb{R}\mathbb{P}^{3}$ are called isotopic if there is an ambient isotopy of $\mathbb{R}\mathbb{P}^{3}$ that maps $L_{0}$ to $L_{1}$. Then their preimages by the covering map $\pi$ are also isotopic in $S^{3}$ since the isotopy is lifted. The main theorem asserts that the converse also holds.
@@
We will prove the main theorem by a similar argument to that in [21]. In Section 2, we will show the main theorem for the hyperbolic links using the Mostow rigidity and the spherical space form conjecture. In Section 3, we will show the main theorem for the links whose complements are Seifert fibered. We will explicitly describe these links, which are generalizations of the torus links. In Section 4, we will show the main theorem for all the links. By the prime and JSJ decompositions of the complement, we are reduced to the above cases and a result for links in the solid torus [21, Theorem 4.5]. Then we need the notion of an outermost JSJ piece as a starting point. Remark that the main theorem holds for the oriented links in $\mathbb{R}\mathbb{P}^{3}$. This can be shown in the same manner by being careful with the orientation.
@@ 2. Hyperbolic links
In this section, we show that a free 2-periodicity of hyperbolic links in $S^{3}$ is unique. For this purpose, we need to use the spherical space form conjecture (now a theorem) as follows. This is a consequence of the geometrization of 3-manifolds and 3-orbifolds and was explicitly shown by Dinkelbach and Leeb [15] using equivariant Ricci flow with surgery.
@@
We consider a condition in a general setting. Let $L_{0}$ and $L_{1}$ be hyperbolic links in $X$. In other words, the complement $M_{i}$ of $L_{i}$ for each $i=0,1$ admits a hyperbolic metric of finite volume. Let $\pi\colon\widetilde{X}\to X$ be a finite regular covering. Suppose that the preimages $\widetilde{L_{0}}=\pi^{-1}(L_{0})$ and $\widetilde{L_{1}}=\pi^{-1}(L_{1})$ are isotopic in $\widetilde{X}$. Then the complement $\widetilde{M_{i}}$ of $\widetilde{L_{i}}$ is a finite cover of $M_{i}$ and is endowed with the lifted hyperbolic metric. The isotopy between $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$ induces a homeomorphism $f\colon\widetilde{M_{0}}\to\widetilde{M_{1}}$, which is isotopic to an isometry $\tilde{\iota}\colon\widetilde{M_{0}}\to\widetilde{M_{1}}$ by the Mostow rigidity. The finite group $G=\pi_{1}(X)/\pi_{1}(\widetilde{X})$ acts on $\widetilde{X}$ by the deck transformations. This action induces free isometric actions $\alpha_{0}$ and $\alpha_{1}$ of $G$ on the hyperbolic 3-manifolds $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$. The isometry $\tilde{\iota}\colon\widetilde{M_{0}}\to\widetilde{M_{1}}$ induces the $G$-action $\tilde{\iota}_{*}(\alpha_{0})$ on $\widetilde{M_{1}}$ from the $G$-action $\alpha_{0}$ on $\widetilde{M_{0}}$.
@@
Suppose that $A\in\mathrm{O}(4)$ acts on $S^{3}$ freely and $A^{2}=I$, where $I\in\mathrm{O}(4)$ is the identity matrix. Since $A^{2}=I$, an eigenvalue of $A$ is 1 or $-1$. Since $A$ acts on $S^{3}$ freely, 1 is not an eigenvalue of $A$. Since $-1$ is the only eigenvalue of $A$ and $A^{2}=I$, a Jordan normal form of $A$ is $-I$. Since $A$ is in the center of $\mathrm{GL}(4,\mathbb{C})$, we have $A=-I$. ∎
@@
In the above setting, let $X=\mathbb{R}\mathbb{P}^{3}$ and $G=\mathbb{Z}/2\mathbb{Z}$. We consider the two subgroups $G_{0}=\tilde{\iota}_{*}(\alpha_{0})(G)$ and $G_{1}=\alpha_{1}(G)$ of the finite group $\operatorname{Isom}(\widetilde{M_{1}})$. Let $\overline{G}$ denote the subgroup of $\operatorname{Isom}(\widetilde{M_{1}})$ generated by $G_{0}$ and $G_{1}$. Since the elements of $G_{0}$ and $G_{1}$ preserves the meridians of $\widetilde{L_{1}}$, so does the elements of $\overline{G}$. Hence the action of the finite group $\overline{G}$ on $\widetilde{M_{1}}$ extends to a smooth action of $S^{3}$. We may assume that this action is isometric by Theorem 2.1. Since $G_{0}$ and $G_{1}$ acts freely and isometrically, the subgroups $G_{0}$ and $G_{1}$ coincide by Lemma 2.4. Lemmas 2.2 and 2.3 imply that $L_{0}$ and $L_{1}$ are isotopic. ∎
@@ 3. Links whose complements are Seifert fibered
For $(p,q)\in\mathbb{Z}^{2}$ and $n\in\{0,1,2\}$, we define the links $T_{S^{3}}(p,q;n)\subset S^{3}$ and $T_{\mathbb{R}\mathbb{P}^{3}}(p,q;n)\subset\mathbb{R}\mathbb{P}^{3}$ as follows. If $(p,q)\neq(0,0)$, let $d=\gcd(p,q)$, $p^{\prime}=p/d$, $q^{\prime}=q/d$, and let $T_{*}(p,q;0)$ for $*=S^{3}$ or $\mathbb{R}\mathbb{P}^{3}$ denote the union of $d$ parallel copies of the simple closed curve on $\Sigma$ representing $p^{\prime}\lambda+q^{\prime}\mu\in H_{1}(\Sigma,\mathbb{Z})$. Conventionally, let $T_{*}(0,0;0)$ denote the empty set. Let $T_{*}(p,q;1)$ denote the union of $T_{*}(p,q;0)$ and the core of $H_{0}$. Let $T_{*}(p,q;2)$ denote the union of $T_{*}(p,q;0)$ and the cores of $H_{0}$ and $H_{1}$. Note that $T_{S^{3}}(p,q;0)$ is the $(p,q)$-torus link.
@@
If two links $T_{S^{3}}(p_{0},q_{0};n_{0})$ and $T_{S^{3}}(p_{1},q_{1};n_{1})$ are isotopic, they are related by combining the relations $(1)$$(4)$.
@@
    If $-p+2q>0$ divides $q$, then $T_{\mathbb{R}\mathbb{P}^{3}}(p,q;1)\cong T_{\mathbb{R}\mathbb{P}^{3}}((-p+2q-1)p/(-p+2q),(-p+2q-1)q/(-p+2q);2)$.
@@
We next prepare notations for Seifert fibrations. For coprime integers $a>0$ and $b$, the $(a,b)$-fibered solid torus is obtained from $D^{2}\times[0,1]$ by identifying $D^{2}\times\{0\}$ and $D^{2}\times\{1\}$ through the $2\pi b/a$-rotation, and its fibers are obtained from the fibers $\{z\}\times[0,1]$ of $D^{2}\times[0,1]$. A $(1,0)$-fibered solid torus is also called a product fibered solid torus. A Seifert fiberation of a 3-manifold is a decomposition into a disjoint union of circles (called fibers) such that each fiber has a tubular neighborhood that forms a fibered solid torus. A fiber is called regular if it has a tubular neighborhood that forms a product fibered solid torus. A fiber is called $(a,b)$-singular if it has a tubular neighborhood that forms an $(a,b)$-fibered solid torus for $a\geq 2$. The following fact is well known.
@@
Let $M$ be a Seifert fibered 3-manifold with boundary consisting of tori. Let $T$ be a component of the boundary of $M$, which is a union of fibers. Suppose that a 3-manifold $M^{\prime}$ is obtained by attaching a solid torus to $M$ along $T$. Then the Seifert fibration of $M$ is uniquely extended to a Seifert fibration of $M^{\prime}$ by a fibered solid torus unless a meridian (i.e. a simple closed curve on $T$ bounding a disk in the solid torus) is isotopic to a fiber in $T$.
@@
We consider the links $T_{*}(p,q;n)$ in relation to Seifert fibrations. For $(p,q)\in\mathbb{Z}^{2}\setminus\{(0,0)\}$, take a Seifert fibration of a tubular neighborhood of the Heegaard torus $\Sigma$ so that a link $T_{*}(p,q;0)$ consists of fibers. This Seifert fibration has no singular fibers. By Lemma 3.2, this can be extended to a Seifert fibration of $S^{3}$ or $\mathbb{R}\mathbb{P}^{3}$ if the slope $p^{\prime}\lambda+q^{\prime}\mu$ is neither the meridian of $H_{0}$ nor $H_{1}$. Then the cores of $H_{0}$ and $H_{1}$ are (in general singular) fibers. Hence a link $T_{*}(p,q;n)$ consists of fibers of a Seifert fibration unless it is isotopic to $T_{*}(0,q;n)$ or $T_{\mathbb{R}\mathbb{P}^{3}}(2q,q;1)$. Note that $T_{S^{3}}(p,0;1)\cong T_{S^{3}}(p+1,0;0)\cong T_{S^{3}}(0,p+1;0)$ for $p\geq 0$. The knot $T_{S^{3}}(0,0;1)\cong T_{S^{3}}(0,1;0)$ is the unknot, and so its complement is a solid torus. The complement of the knot $T_{\mathbb{R}\mathbb{P}^{3}}(0,0;1)\cong T_{\mathbb{R}\mathbb{P}^{3}}(1,q;0)$ is also a solid torus. The other $T_{*}(0,q;0)$ and $T_{\mathbb{R}\mathbb{P}^{3}}(2q,q;1)$ are not irreducible, and so they are not Seifert fibered. Note that a Seifert fibered 3-manifold is irreducible unless it is $S^{1}\times S^{2}$, the twisted $S^{2}$-bundle over $S^{1}$, or $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ [19, Proposition 1.12]. The link $T_{S^{3}}(0,q;1)\cong T_{S^{3}}(0,q-1;2)$ for $q\geq 2$ is called a “key-chain link” in [9]. The following homeomorphisms hold.
@@
By Proposition 3.4, we may assume that $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};n_{0})$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{1},q_{1};n_{1})$. By Lemma 3.5, we have $\widetilde{L_{0}}=T_{S^{3}}(p_{0},-p_{0}+2q_{0};n_{0})$ and $\widetilde{L_{1}}=T_{S^{3}}(p_{1},-p_{1}+2q_{1};n_{1})$, which are isotopic. We need to check the relations (1)–(4) in Proposition 3.1. In general cases for the relations (1) and (2), $n_{0}=n_{1}$ and $(p_{0},-p_{0}+2q_{0})$ is equal to ±$\pm(p_{1},-p_{1}+2q_{1})$ or ±$\pm(-p_{1}+2q_{1},p_{1})$. Then $(p_{0},q_{0})$ is equal to ±$\pm(p_{1},q_{1})$ or ±$\pm(-p_{1}+2q_{1},q_{1})$. Hence $L_{0}$ and $L_{1}$ are isotopic.
@@
  • Suppose that $p_{0}>0$ divides $-p_{0}+2q_{0}$, $n_{0}=0$, $n_{1}=1$, $p_{1}=p_{0}-1$, and $-p_{1}+2q_{1}=(p_{0}-1)(-p_{0}+2q_{0})/p_{0}$. Note that $T_{S^{3}}(p_{0},-p_{0}+2q_{0};1)\cong T_{S^{3}}(p_{0}-1,(p_{0}-1)(-p_{0}+2q_{0})/p_{0};0)$. Then $q_{1}=(p_{0}-1)q_{0}/p_{0}$ and $p_{0}$ divides $q_{0}$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};0)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0}-1,(p_{0}-1)q_{0}/p_{0};1)$ are isotopic.
  • Suppose that $-p_{0}+2q_{0}>0$ divides $p_{0}$, $n_{0}=1$, $n_{1}=2$, $p_{1}=(-p_{0}+2q_{0}-1)p_{0}/(-p_{0}+2q_{0})$, and $-p_{1}+2q_{1}=-p_{0}+2q_{0}-1$. Note that $T_{S^{3}}(p_{0},-p_{0}+2q_{0};2)\cong T_{S^{3}}((-p_{0}+2q_{0}-1)p_{0}/(-p_{0}+2q_{0}),-p_{0}+2q_{0}-1;1)$. Then $q_{1}=(-p_{0}+2q_{0}-1)q_{0}/(-p_{0}+2q_{0})$ and $-p_{0}+2q_{0}$ divides $q_{0}$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};1)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}((-p_{0}+2q_{0}-1)p_{0}/(-p_{0}+2q_{0}),(-p_{0}+2q_{0}-1)q_{0}/(-p_{0}+2q_{0});2)$ are isotopic.
  • Suppose that $p_{0}=\pm(-p_{0}+2q_{0})>0$, $n_{0}=0$, $n_{1}=2$, $p_{1}=p_{0}-2$, and $-p_{1}+2q_{1}=\pm(p_{0}-2)$. Note that $T_{S^{3}}(p_{0},-p_{0}+2q_{0};0)=T_{S^{3}}(p_{0},\pm p_{0};2)\cong T_{S^{3}}(p_{0}-2,\pm(p_{0}-2);2)$. If $p_{0}=-p_{0}+2q_{0}$, then $q_{0}=p_{0}$ and $q_{1}=p_{0}-2$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};0)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0}-2,p_{0}-2;2)$ are isotopic. If $p_{0}=-(-p_{0}+2q_{0})$, then $q_{0}=q_{1}=0$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};0)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0}-2,0;2)$ are isotopic.
We finally remark that the relations (1)–(4) for the isotopy classes of $T_{\mathbb{R}\mathbb{P}^{3}}(p,q;n)$ are sufficient, for otherwise more preimages in $S^{3}$ would be identified.
@@
If two links $T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};n_{0})$ and $T_{\mathbb{R}\mathbb{P}^{3}}(p_{1},q_{1};n_{1})$ are isotopic, they are related by combining the relations $(1)$$(4)$.

### v2

@@ 1. Introduction
Links in the real projective 3-space $\mathbb{R}\mathbb{P}^{3}$ (also known as projective links) have been investigated in various ways, often using diagrams, as in [16, 28, 29]. More generally, links in lens spaces have been considered [1, 10]. For a link $L$ in a lens space, the preimage (also called the lift) of $L$ by the universal covering map of the lens space is a freely periodic link in the 3-sphere $S^{3}$ according to Hartley [18]. Free periodicities of links in $S^{3}$ are quite restricted, especially for knots. For instance, Sakuma [33] and Boileau and Flapan [2] independently proved that a free $n$-periodicity of a non-trivial prime knot in $S^{3}$ is unique for each $n\geq 2$. More precisely, for a non-trivial prime knot $K$ in $S^{3}$ and the group $\operatorname{Diff}^{*}(S^{3},K)$ consisting of the diffeomorphisms of the pair $(S^{3},K)$ which preserve the orientations of $S^{3}$ and $K$, if $f,g\in\operatorname{Diff}^{*}(S^{3},K)$ have no fixed points and satisfy $f^{n}=g^{n}=\mathrm{id}$, then $f$ is conjugate in $\operatorname{Diff}^{*}(S^{3},K)$ to a power of $g$. Equivalently, if $K_{0}\subset L(n,q_{0})$ and $K_{1}\subset L(n,q_{1})$ are knots in lens spaces, and their premiages in $S^{3}$ are isotopic knots, then the pairs $(L(n,q_{0}),K_{0})$ and $(L(n,q_{1}),K_{1})$ are diffeomorphic. Several constraints for freely periodic knots in $S^{3}$ are known [5, 6, 12, 13, 18, 31]. Compatibility of free periodicities and alternating diagrams has also been considered [7, 14]. Furthermore, the finite group actions on knots in $S^{3}$ were classified in [4].
In this paper, we consider links in $\mathbb{R}\mathbb{P}^{3}$ and their preimages in $S^{3}$. Every link is assumed to be tame. Let $\pi\colon S^{3}\to\mathbb{R}\mathbb{P}^{3}$ denote the universal covering map, which is a double cover. Two links $L_{0}$ and $L_{1}$ in $\mathbb{R}\mathbb{P}^{3}$ are called isotopic if there is an ambient isotopy of $\mathbb{R}\mathbb{P}^{3}$ that maps $L_{0}$ to $L_{1}$. Then their preimages by the covering map $\pi$ are also isotopic in $S^{3}$ since the isotopy is lifted. The main theorem asserts that the converse also holds.
@@
We will prove the main theorem by a similar argument to that in [24]. In Section 2, we will show the main theorem for the hyperbolic links using the Mostow rigidity theorem and the resolution of the spherical space form conjecture. In Section 3, we will show the main theorem for the links whose complements are Seifert fibered. We will explicitly describe these links, which are generalizations of the torus links. In Section 4, we will show the main theorem for all links. By the prime and JSJ decompositions of the complement, we reduce to the above cases and a result for links in the solid torus [24, Theorem 4.5]. Then we need the notion of an outermost JSJ piece as a starting point. We remark that the main theorem holds for the oriented links in $\mathbb{R}\mathbb{P}^{3}$. This can be shown in the same manner by being careful with the orientation.
@@ 2. Hyperbolic links
In this section, we show that a free 2-periodicity of hyperbolic links in $S^{3}$ is unique. For this purpose, we need to use the now proven spherical space form conjecture as follows. This is a consequence of the geometrization of 3-manifolds and 3-orbifolds and was explicitly shown by Dinkelbach and Leeb [15] using equivariant Ricci flow with surgery.
@@
[…]
@@
Suppose that $A\in\mathrm{O}(4)$ acts on $S^{3}$ freely and $A^{2}=I$, where $I\in\mathrm{O}(4)$ is the identity matrix. Since $A^{2}=I$, an eigenvalue of $A$ is 1 or 1-1. Since $A$ acts on $S^{3}$ freely, 1 is not an eigenvalue of $A$. Since 1-1 is the only eigenvalue of $A$ and $A^{2}=I$, a Jordan normal form of $A$ is I-I. Since $A$ is in the center of $\mathrm{GL}(4,\mathbb{C})$, we have A=IA=-I. ∎
@@
[…]
Remark 2.6.
In the above proof, we obtain that the finite group $\overline{G}$ is actually $\mathbb{Z}/2\mathbb{Z}$ after we obtain a (not necessarily free) isometric action of $\overline{G}$ on $S^{3}$. Hence we need the resolution of the spherical space form conjecture in full generality, as well as the case of a free involution shown by Livesay [25].
@@ 3. Links whose complements are Seifert fibered
For $(p,q)\in\mathbb{Z}^{2}$ and $n\in\{0,1,2\}$, we define the links $T_{S^{3}}(p,q;n)\subset S^{3}$ and $T_{\mathbb{R}\mathbb{P}^{3}}(p,q;n)\subset\mathbb{R}\mathbb{P}^{3}$ as follows. If $(p,q)\neq(0,0)$, let $d=\gcd(p,q)$, $p^{\prime}=p/d$, $q^{\prime}=q/d$, and let $T_{*}(p,q;0)$ for $S^{3}$ or $\mathbb{R}\mathbb{P}^{3}$ denote the union of $d$ parallel copies of the simple closed curve on $\Sigma$ representing $p^{\prime}\lambda+q^{\prime}\mu\in H_{1}(\Sigma,\mathbb{Z})$. Conventionally, let $T_{*}(0,0;0)$ denote the empty set. Let $T_{*}(p,q;1)$ denote the union of $T_{*}(p,q;0)$ and the core of $H_{0}$. Let $T_{*}(p,q;2)$ denote the union of $T_{*}(p,q;0)$ and the cores of $H_{0}$ and $H_{1}$. Note that $T_{S^{3}}(p,q;0)$ is the $(p,q)$-torus link.
@@
If two links $T_{S^{3}}(p_{0},q_{0};n_{0})$ and $T_{S^{3}}(p_{1},q_{1};n_{1})$ are isotopic, they are related by combining the relations (1)(1)$(4)$.
@@
    If p+2q>0-p+2q>0 divides $q$, then $T_{\mathbb{R}\mathbb{P}^{3}}(p,q;1)\cong T_{\mathbb{R}\mathbb{P}^{3}}((-p+2q-1)p/(-p+2q),(-p+2q-1)q/(-p+2q);2)$.
@@
We next prepare notations for Seifert fibrations. For coprime integers $a>0$ and $b$, the $(a,b)$-fibered solid torus is obtained from $D^{2}\times[0,1]$ by identifying $D^{2}\times\{0\}$ and $D^{2}\times\{1\}$ through the $2\pi b/a$-rotation, and its fibers are obtained from the fibers $\{z\}\times[0,1]$ of $D^{2}\times[0,1]$. A $(1,0)$-fibered solid torus is also called a product fibered solid torus. A Seifert fibration of a 3-manifold is a decomposition into a disjoint union of circles (called fibers) such that each fiber has a tubular neighborhood that forms a fibered solid torus. A fiber is called regular if it has a tubular neighborhood that forms a product fibered solid torus. A fiber is called $(a,b)$-singular if it has a tubular neighborhood that forms an $(a,b)$-fibered solid torus for $a\geq 2$. The following fact is well known.
@@
Let $M$ be a Seifert fibered 3-manifold with boundary consisting of tori. Let $T$ be a component of the boundary of $M$, which is a union of fibers. Suppose that a 3-manifold $M^{\prime}$ is obtained by attaching a solid torus to $M$ along $T$. Then the Seifert fibration of $M$ is uniquely extended to a Seifert fibration of $M^{\prime}$ by a fibered solid torus unless a meridian (i.e., a simple closed curve on $T$ bounding a disk in the solid torus) is isotopic to a fiber in $T$.
@@
We now consider the links $T_{*}(p,q;n)$ in relation to Seifert fibrations. For $(p,q)\in\mathbb{Z}^{2}\setminus\{(0,0)\}$, take a Seifert fibration of a tubular neighborhood of the Heegaard torus $\Sigma$ so that a link $T_{*}(p,q;0)$ consists of fibers. This Seifert fibration has no singular fibers. By Lemma 3.2, this can be extended to a Seifert fibration of $S^{3}$ or $\mathbb{R}\mathbb{P}^{3}$ if the slope $p^{\prime}\lambda+q^{\prime}\mu$ is neither the meridian of $H_{0}$ nor $H_{1}$. Then the cores of $H_{0}$ and $H_{1}$ are (in general singular) fibers. Hence a link $T_{*}(p,q;n)$ consists of fibers of a Seifert fibration unless it is isotopic to $T_{*}(0,q;n)$ or $T_{\mathbb{R}\mathbb{P}^{3}}(2q,q;1)$. Note that $T_{S^{3}}(p,0;1)\cong T_{S^{3}}(p+1,0;0)\cong T_{S^{3}}(0,p+1;0)$ for $p\geq 0$. The knot $T_{S^{3}}(0,0;1)\cong T_{S^{3}}(0,1;0)$ is the unknot, and so its complement is a solid torus. The complement of the knot $T_{\mathbb{R}\mathbb{P}^{3}}(0,0;1)\cong T_{\mathbb{R}\mathbb{P}^{3}}(1,q;0)$ is also a solid torus. The other links $T_{*}(0,q;0)$ ($q\geq 2$) and $T_{\mathbb{R}\mathbb{P}^{3}}(2q,q;1)$ ($q\geq 1$) are not irreducible, and so they are not Seifert fibered. Note that a Seifert fibered 3-manifold is irreducible unless it is $S^{1}\times S^{2}$, the twisted $S^{2}$-bundle over $S^{1}$, or $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ [19, Proposition 1.12]. The link $T_{S^{3}}(0,q;1)\cong T_{S^{3}}(0,q-1;2)$ for $q\geq 2$ is called a “key-chain link” in [9]. The following homeomorphisms hold.
@@
By Proposition 3.4, we may assume that $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};n_{0})$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{1},q_{1};n_{1})$. By Lemma 3.5, we have $\widetilde{L_{0}}=T_{S^{3}}(p_{0},-p_{0}+2q_{0};n_{0})$ and $\widetilde{L_{1}}=T_{S^{3}}(p_{1},-p_{1}+2q_{1};n_{1})$, which are isotopic. We need to check the relations (1)–(4) in Proposition 3.1. In general cases for the relations (1) and (2), $n_{0}=n_{1}$ and $(p_{0},-p_{0}+2q_{0})$ is equal to $\pm(p_{1},-p_{1}+2q_{1})$ or $\pm(-p_{1}+2q_{1},p_{1})$. Then $(p_{0},q_{0})$ is equal to $\pm(p_{1},q_{1})$ or $\pm(-p_{1}+2q_{1},q_{1})$. Hence $L_{0}$ and $L_{1}$ are isotopic.
@@
  • Suppose that $p_{0}>0$ divides $p_{0}+2q_{0}$, $n_{0}=0$, $n_{1}=1$, $p_{1}=p_{0}-1$, and $p_{1}+2q_{1}=(p_{0}-1)(-p_{0}+2q_{0})/p_{0}$. Note that $T_{S^{3}}(p_{0},-p_{0}+2q_{0};1)\cong T_{S^{3}}(p_{0}-1,(p_{0}-1)(-p_{0}+2q_{0})/p_{0};0)$. Then $q_{1}=(p_{0}-1)q_{0}/p_{0}$ and $p_{0}$ divides $q_{0}$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};0)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0}-1,(p_{0}-1)q_{0}/p_{0};1)$ are isotopic.
  • Suppose that $p_{0}+2q_{0}>0$ divides $p_{0}$, $n_{0}=1$, $n_{1}=2$, $p_{1}=(-p_{0}+2q_{0}-1)p_{0}/(-p_{0}+2q_{0})$, and $p_{1}+2q_{1}=-p_{0}+2q_{0}-1$. Note that $T_{S^{3}}(p_{0},-p_{0}+2q_{0};2)\cong T_{S^{3}}((-p_{0}+2q_{0}-1)p_{0}/(-p_{0}+2q_{0}),-p_{0}+2q_{0}-1;1)$. Then $q_{1}=(-p_{0}+2q_{0}-1)q_{0}/(-p_{0}+2q_{0})$ and $p_{0}+2q_{0}$ divides $q_{0}$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};1)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}((-p_{0}+2q_{0}-1)p_{0}/(-p_{0}+2q_{0}),(-p_{0}+2q_{0}-1)q_{0}/(-p_{0}+2q_{0});2)$ are isotopic.
  • Suppose that $p_{0}=\pm(-p_{0}+2q_{0})>0$, $n_{0}=0$, $n_{1}=2$, $p_{1}=p_{0}-2$, and $p_{1}+2q_{1}=\pm(p_{0}-2)$. Note that $T_{S^{3}}(p_{0},-p_{0}+2q_{0};0)=T_{S^{3}}(p_{0},\pm p_{0};2)\cong T_{S^{3}}(p_{0}-2,\pm(p_{0}-2);2)$. If $p_{0}=-p_{0}+2q_{0}$, then $q_{0}=p_{0}$ and $q_{1}=p_{0}-2$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};0)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0}-2,p_{0}-2;2)$ are isotopic. If $p_{0}=-(-p_{0}+2q_{0})$, then $q_{0}=q_{1}=0$. Hence $L_{0}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};0)$ and $L_{1}=T_{\mathbb{R}\mathbb{P}^{3}}(p_{0}-2,0;2)$ are isotopic.
We finally remark that the relations (1)–(4) for the isotopy classes of $T_{\mathbb{R}\mathbb{P}^{3}}(p,q;n)$ are sufficient, for otherwise more preimages in $S^{3}$ would be identified. Hence:
@@
If two links $T_{\mathbb{R}\mathbb{P}^{3}}(p_{0},q_{0};n_{0})$ and $T_{\mathbb{R}\mathbb{P}^{3}}(p_{1},q_{1};n_{1})$ are isotopic, they are related by combining the relations (1)(1)$(4)$.